OpenStudy (anonymous):

differentiating (x-y)^2=x+y-1

OpenStudy (anonymous):

wrt x yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so you start with \[(x-y)^2=x+y+1\] \[2(x-y)(1+y')=1+y'\]

OpenStudy (anonymous):

is that much ok?

OpenStudy (anonymous):

in implicit differentiation you are thinking that \[y=f(x)\] so \[y'=f'(x)\]

OpenStudy (anonymous):

I don't know how to get to the next step: 2 (x-y) - 2 (x-y) y' = 1 + y' are you using the chain rule or something to get to it?

OpenStudy (anonymous):

oh if the first step is ok the we are on the way

OpenStudy (anonymous):

now it is just annoying algebra to solve for y'

OpenStudy (anonymous):

er. i mean how to get to the step after the one you posted and to the step that i posted.

OpenStudy (anonymous):

i guess you have to multiply out on the left, collect terms, put everything with y' on one side and everything else on the other, factor and divide

OpenStudy (anonymous):

that is the answer that the solution manual gives, but it is lacking a coherent explanation

OpenStudy (anonymous):

the fist step is the only "calculus " step. it is algebra from here on in

OpenStudy (anonymous):

we can work it out step by step if you like

OpenStudy (anonymous):

yes please :)

OpenStudy (anonymous):

ok but first is this ok? \[2(x-y)(1+y')=1+y'\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if not we should work that out first

OpenStudy (anonymous):

that is actually where i got stuck on

OpenStudy (anonymous):

wait you got stuck on first step? then lets do it slowly

OpenStudy (anonymous):

first of all right hand side is just \[x+y+1\] and we are thinking that \[y=f(x)\] even though we don't know what \[f(x)\] is i.e. we have not "solved" for it

OpenStudy (anonymous):

that is why it is called "implicit differentiation because we are thinking of y "implicitly" as a function of x

OpenStudy (anonymous):

so if we write the right hand side as \[x+f(x)+1\] and take derivatives we get \[1+f'(x)\] because the derivative of x is 1, the derivative of f(x) is just f'(x) and the derivative of 1 is 0

OpenStudy (anonymous):

that is why the right hand side is \[1+y'\]

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok now to the left hand side \[(x-y)^2\]

OpenStudy (anonymous):

which i will write as \[(x-f(x))^2\]

OpenStudy (anonymous):

by the chain rule the derivative of something squared is twice something times the derivative of something. so i write \[2(x-f(x)) \times \frac{d}{dx}(x-f(x))\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

the the derivative of x is 1, the derivative of f(x) is f'(x) so we get \[2(x-f(x))(1-f'(x))\] or \[2(x-y)(1-y')\]

OpenStudy (anonymous):

so we are back to \[(2(x-y)(1-y')=1+y'\]and now our only job is to solve this equation for \[y'\]

OpenStudy (anonymous):

ugly algebra but we can work it out

OpenStudy (anonymous):

\[2(x-xy'-y+yy')=1+y'\] \[2x-2xy'-2y+2yy'=1+y'\] \[2x-2y-1=y'+2xy'-2yy'\]

OpenStudy (anonymous):

\[2x-2y-1=y'(1+2x-2y)\] \[\frac{2x-2y-1}{1+2x-2y}=y'\]

OpenStudy (anonymous):

ah shucks. i completely failed to realize that you could foil.

OpenStudy (anonymous):

there might have been a snappier way to solve but i did not see it. also i could have made an algebra mistake so check my work

OpenStudy (anonymous):

curses foiled again

OpenStudy (anonymous):

haha.. thanks for taking the time to explain it!

OpenStudy (anonymous):

really you should check my algebra because i just typed it out and didn't do it with pencil and paper

OpenStudy (anonymous):

i am going to in a second. just going to do it on my notebook :P

OpenStudy (anonymous):

hope it helped. good luck. are you done or do you have another one?

OpenStudy (anonymous):

atm no. though i might skip to related rates and then yes.

OpenStudy (anonymous):

related rates! must have just passed the chair rule part of the course. sliding ladders and oil spills, planes flying overhead and the guy walking away from the lamp. have fun!